OpenStudy (anonymous):

A local gym charges nonmembers $10 per hour to use the tennis courts. Members pay a yearly fee of $300 and $4 per hour for using the tennis courts. Write an equation to find how many hours, h, you must use the tennis courts to justify becoming a member.

OpenStudy (anonymous):

PLEASE HELP

jimthompson5910 (jim_thompson5910):

nonmembers: Total Cost = 10h Members: Total Cost = 300 + 4h Equate the two to get 10h = 300 + 4h then solve for h to get your answer

OpenStudy (anonymous):

I DONT NEED TO FIND H I NEED A EQUATION

jimthompson5910 (jim_thompson5910):

your equation is 10h = 300 + 4h, and please try not to shout

jimthompson5910 (jim_thompson5910):

you would solve that for h to get the number of hours where the two costs are the same anything after that is when the membership is cheaper

OpenStudy (anonymous):

sorry it was on caps

jimthompson5910 (jim_thompson5910):

thats ok

OpenStudy (anonymous):

thanks though big help

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

sure a few more

OpenStudy (anonymous):

ok cool. Rewrite the equation so that y is a function of x: 3x + 5y = 15

jimthompson5910 (jim_thompson5910):

you want to solve for y

jimthompson5910 (jim_thompson5910):

so you must move that 3x over how do you do this?

OpenStudy (anonymous):

idk

jimthompson5910 (jim_thompson5910):

3x + 5y = 15 is the same as 5y + 3x = 15 so you want to undo the +3x

jimthompson5910 (jim_thompson5910):

how do you undo +3x ??

OpenStudy (anonymous):

i d k how to these problems at all

jimthompson5910 (jim_thompson5910):

what's the opposite of addition

OpenStudy (anonymous):

subtraction.

jimthompson5910 (jim_thompson5910):

so you have to subtract 3x from both sides

jimthompson5910 (jim_thompson5910):

5y + 3x = 15 5y + 3x-3x = 15-3x 5y + 0x = 15-3x 5y = 15-3x 5y = -3x + 15 Then what's next?

OpenStudy (anonymous):

I don't get it?

jimthompson5910 (jim_thompson5910):

you have 5y or 5 times y how do you undo multiplication?